# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v2

@@ 6.1. Cut-and-paste procedure: nonorientable setting
The nonorientable compact four-manifold $N_{p}$ of the second step of our procedure is constructed as follows; cf. Section 2.1. Denote the total space of the nonorientable three-sphere bundle over the circle by $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ and let $\alpha\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=\mathbb{Z}$. Let $p$ be an odd integer and denote by $\alpha_{p}\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ the orientation-reversing simple loop representing the class $[\alpha]^{p}\in H_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z})=\mathbb{Z}$. Its tubular neighborhood is diffeomorphic to the total space of the nonorientable three-disk bundle over the circle $\nu(\alpha_{p})=D^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$.
@@
For any odd $p\in\mathbb{Z}_{>0}$, the compact nonorientable four-manifold with boundary $N_{p}$ defined in (6.1) is double covered by the orientable four-manifold $S_{p}$ (2.1):
$S_{p}\xrightarrow{2:1}N_{p}.$
@@
$\psi:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
given by
$(t,x)\mapsto(t+1,g_{p}(x)),$
where $g_{p}$ is the restriction of an orientation reversing automorphism of $S^{3}$ that cyclically permutes $B_{1},\dots,B_{p}$. The conclusion follows from the existence of a diffeomorphism
$S_{p}\rightarrow(\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i}))/\sim_{\psi^{2}}.$
@@
We can view $S^{3}$ by means of its genus one Heegaard splitting (2.3) and assume that the map $g_{p}$ in the proof of Lemma 23 is the composition of the map $f_{p}$ of Remark 3 and a reflection $\rho$ with respect to a two-sphere that intersects the first solid torus of (2.3) in $S^{1}\times D^{1}$ and the second one in the disjoint union of two $D^{2}$-fibers.
@@
Theorem A follows from Theorem 4 and Theorem 25.
[Proof]Kim-Raymond [24] (cf. [30]) showed that the diffeotopy group of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ is a quotient of the one of $S^{1}\times S^{2}$. This is isomorphic to $\mathbb{Z}/2\times\mathbb{Z}/2\times\mathbb{Z}/2$ and a set of generators is given by the Gluck twist (2.4) and the product maps $Id_{S^{1}}\times a$ and $r\times Id_{S^{2}}$, where $a$ is the antipodal involution of $S^{2}$ and $r$ is a orientation-reversing automorphism of $S^{1}$. We prove the extension result for the map $f:S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}\rightarrow S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ given by the Gluck twist and leave the cases of the other generators of the diffeotopy group of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ to the interested reader.
Without loss of generality, we can work in the context of Remark 24. Consider the map
$G:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow S^{3}\setminus\sqcup_{i=1}^{p}B_{i}$
$(t,x)\mapsto r_{t}(x)$
where $r_{t}$ is a rotation of $S^{3}$ of angle $\pi t$, restricting to a rotation along the $D^{2}$-factor of $D^{2}\times S^{1}$ in (2.3). One can check that the smooth automorphism
$T:\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\sqcup_{i=1}^{p}B_{i})$
$(t,x)\mapsto(t,G(t,x))$
is $\psi$-equivariant (where $\psi$ is the map generating the deck transformations of $N_{p}$ in Lemma 23) and, therefore, it induces a smooth automorphism
$t_{p}:N_{p}\rightarrow N_{p}.$
In particular, the restriction of $t_{p}$ to the boundary is obtained by composing $p$ copies of the automorphism $t$. This yields a map isotopic to $t$, since $p$ was assumed to be odd.
@@
The simplicity of the construction makes it straight-forward to compute several basic invariants of the new four-manifold (6.2) as we now sample; see [22] for details on these invariants.
@@
    If there is a $\operatorname{Pin}^{\mathbb{C}}$-structure $(Y,\phi^{\mathbb{C}}_{Y})$, then there is a $\operatorname{Pin}^{\mathbb{C}}$-structure $(Y(p),\phi^{\mathbb{C}}_{Y(p)})$ and the Arf invariant is $\operatorname{arf}(Y,\phi^{\mathbb{C}}_{Y})=\operatorname{arf}(Y(p),\phi^{\mathbb{C}}_{Y(p)})$.
  4. (4)
@@ 6.2. Nonorientable raw material and its properties
$\bullet$ Gompf [19]. The universal covering $\widetilde{Q}$ is diffeomorphic to $S^{4}$.
@@
$Q\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\rightarrow{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}.$
@@
where $(d,x)\sim(-d,\mathbb{A}(x))$ for every $d\in D^{2}$ and $x\in S^{2}$ and $\mathbb{A}:S^{2}\rightarrow S^{2}$ is the antipodal involution. The analysis done by Akbulut [1, Section 1] on the four-manifold $Q$ built by Cappell-Shaneson [10] reveals that the compact four-manifold $Z\mathrel{\mathop{:}}=Q\setminus\nu(\alpha_{Q})$ (where $\alpha_{Q}\subset Q$ is a simple loop that generates $\pi_{1}(Q)=\mathbb{Z}/2$) is homeomorphic, but not diffeomorphic to $D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2}$. The second clause of Theorem 28 encapsules work of Gompf, where he showed that the orientation two-covering of $Z$ is diffeomorphic to $D^{2}\times S^{2}$. The third clause of Theorem 28 follows from the property $Z\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}=(D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2})\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}$ [1, Section 3].
@@ 6.3. A spectral invariant as a diffeomorphism obstruction
$\bullet$ Two closed smooth nonorientable $\operatorname{Pin}^{+}$- four-manifolds $(Y,\phi_{Y})$ and $(Y^{\prime},\phi_{Y^{\prime}})$ with $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2=H^{1}(Y^{\prime};\mathbb{Z}/2)$ are stably diffeomorphic if and only if their Euler characteristics agree and $\eta(Y,\phi_{Y})=\eta(Y^{\prime},\phi_{Y^{\prime}})\mod 2\mathbb{Z}$.
@@ 6.4. $\operatorname{Pin}^{+}$-cobordism classes and a computational tool
$\bullet$ Kirby-Taylor [26, Theorem 5.2]. There is a group isomorphism
@@
$[({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})]=\pm 1\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
@@
$[(Q,\pm\phi_{Q})]=\pm 9\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
@@
Let $Y$ be a closed smooth four-manifold with first cohomology group $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2$ and that admits a $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$. Let $\alpha_{Y}\subset Y$ be an orientation-reversing simple loop such that $[\alpha_{Y}]\neq 0\in H_{1}(Y;\mathbb{Z})$ and build
@@
[Proof]We first assemble the $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$ by gluing together $\operatorname{Pin}^{+}$-structures $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ and $(N_{p},\phi_{N_{p}})$ that induce the same $\operatorname{Pin}^{+}$-structure on their common boundary; see (6.3). The $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ induces a $\operatorname{Pin}^{+}$-structure $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ as a codimension zero submanifold [26, Construction 1.5], which restricts to a $\operatorname{Pin}^{+}$-structure
@@
there are two choices of such structures [35, Corollary 6.4] and there is a one-to-one correspondence between the initial $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ and the $\operatorname{Pin}^{+}$-structure (6.6) (see [25, Proposition 1 Chapter IV]). We now argue that there is a $\operatorname{Pin}^{+}$-structure $(N_{p},\phi_{N_{p}})$ that induces the $\operatorname{Pin}^{+}$-structure (6.6) on $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Since $w_{2}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=0$ and $H^{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z}/2)=\mathbb{Z}/2$, there are two $\operatorname{Pin}^{+}$-structures on $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ [35, Corollary 6.4]. The hypothesis $p\neq 0\mod 2$ allows us to pick the $\operatorname{Pin}^{+}$-structure $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ that induces $(N_{p},\phi_{N_{p}})$ such that the induced $\operatorname{Pin}^{+}$-structure $(\partial N_{p},\phi_{N_{p}}|_{\partial})$ matches (6.6). We conclude that there is a $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$. Next, we put together the compact five-manifold $V$ by gluing the product cobordisms
@@
$[(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\pm\phi)]=0\in\Omega^{\operatorname{Pin}^{+}}_{4},$

### v3

@@ 6.1. Cut-and-paste procedure: nonorientable setting
The nonorientable compact four-manifold $N_{p}$ of the second step of our procedure is constructed as follows; cf. Section 2.1. Denote the total space of the nonorientable three-sphere bundle over the circle by $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ and let $\alpha\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ be a simple loop whose homotopy class generates the infinite cyclic group $\pi_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=\mathbb{Z}$. Let $p$ be an odd integer and denote by $\alpha_{p}\subset S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ the orientation-reversing simple loop representing the class $\alpha]^{p}\in H_{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z})=\mathbb{Z}$. Its tubular neighborhood is diffeomorphic to the total space of the nonorientable three-disk bundle over the circle $\nu(\alpha_{p})=D^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$.
@@
For any odd $p\in\mathbb{Z}_{>0}$, the compact nonorientable four-manifold with boundary $N_{p}$ defined in (6.1) is double-covered
$S_{p}\xrightarrow{2:1}N_{p}$
by the orientable four-manifold $S_{p}$ defined in (2.1).
@@
(6.2) $\psi:\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$
defined by
(6.3) $(t,x)\mapsto(t+1,g_{p}(x)),$
where $g_{p}$ is the restriction of an orientation-reversing self-diffeomorphism $S^{3}\rightarrow S^{3}$ that cyclically permutes the three-balls $\{B_{1},\dots,B_{p}\}$. The conclusion of the lemma follows from the existence of a diffeomorphism
(6.4) $S_{p}\rightarrow(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\psi^{2}}.$
@@
Thinking of $S^{3}$ in terms of its genus one Heegaard splitting given in (2.5), we can assume that the map $g_{p}$ in the proof of Lemma 27 is the composition of the map $f_{p}:(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$ of Remark 3 and a reflection with respect to a two-sphere that intersects the solid torus to the left of (2.5) in $S^{1}\times D^{1}$ and the solid torus to the right one the disjoint union of two $D^{2}$-fibers.
@@
Theorem A follows from Theorem 4 and Theorem 29. We denote by
(6.5) $\varphi:S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}\rightarrow S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$
the diffeomorphism given by $\varphi(\theta,x)=(\theta,r_{\theta}(x))$, where $r_{\theta}:S^{2}\rightarrow S^{2}$ is the rotation around a fixed axis of the two-sphere fiber about an angle $\theta$ in the circle base of $S^{2}\widetilde{\times}S^{1}$.
[Proof]The following argument is similar to the one used to prove Theorem 4. Kim-Raymond [26] showed that the diffeotopy group of $S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ is isomorphic to $\mathbb{Z}/2\times\mathbb{Z}/2$, and that it is generated by a reflection and by the diffeomorphism (6.5). We prove the extension result for the latter self-diffeomorphisms since the case of the reflection is immediate. Without loss of generality, we work in the context of Remark 28. Consider the map
(6.6) $G:\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}$
defined as
(6.7) $(t,x)\mapsto r_{t}(x),$
where $r_{t}:S^{3}\rightarrow S^{3}$ is a rotation of an angle of $\pi t$ that restricts to a rotation along the $D^{2}$-factor of the solid torus $D^{2}\times S^{1}$ of the Heegaard decomposition of $S^{3}$ described in Remark 3. Let $\psi$ be the map defined in (6.2) that generates the deck transformations of $N_{p}$ in Lemma 27. The diffeomorphism
(6.8) $T:\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$
defined by
(6.9) $(t,x)\mapsto(t,G(t,x))$
is $\psi$-equivariant and, therefore, it induces a self-diffeomorphism
(6.10) $F_{p}:N_{p}\rightarrow N_{p}.$
To conclude the proof, it is enough to observe that the map
(6.11) $T_{p}:(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})/\sim_{\psi^{2}}\rightarrow(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})/\sim_{\psi^{2}}\cong S_{p}$
induced by the diffeomorphism (6.8) on the double cover $S_{p}\rightarrow N_{p}$ (see (6.2) and Lemma 27) restricts to a Gluck twist on the boundary $\partial S_{p}$. Notice that $T_{p}$ is the lift of $F_{p}$ to $S_{p}$. The fact that such restriction of (6.11) is indeed a Gluck twist follows from the existence of a commutative diagram
$(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\phi}$$(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\phi}$$(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\psi^{2}}$$(\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i}))/\sim_{\psi^{2}},$$\scriptstyle{\Phi_{p}}$$\scriptstyle{m_{2}}$$\scriptstyle{m_{2}}$$\scriptstyle{T_{p}}$
where the horizontal diffeomorphism $\Phi_{p}$ is defined in the proof of Theorem 4, the map $\phi$ is defined in (2.3), and
(6.12) $m_{2}\colon\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})\rightarrow\mathbb{R}\times(S^{3}\setminus\overset{p}{\underset{i=1}{\sqcup}}B_{i})$
is given by $m_{2}(t,x)=(2t,x)$. This concludes the proof of Theorem 29. For the convenience of the reader, we mention that another way to argue is to notice that the restriction of the diffeomorphism (6.10) to the boundary $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$ is the composition of $p$ times the diffeomorphism (6.5). Such composition is isotopic to (6.5) given that $p$ was assumed to be odd by the result of Kim-Raymond [26].
@@
The simplicity of the construction makes it straight-forward to compute several basic invariants of the new four-manifold (6.13) as we now sample.
@@ 6.2. Nonorientable raw material and its properties
$\bullet$ Gompf [20]. The universal cover $\widetilde{Q}$ is diffeomorphic to $S^{4}$.
@@
$Q\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\rightarrow{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}.$
@@
where $(d,x)\sim(-d,\mathbb{A}(x))$ for every $d\in D^{2}$ and $x\in S^{2}$ and $\mathbb{A}:S^{2}\rightarrow S^{2}$ is the antipodal involution. The analysis done by Akbulut [1, Section 1] on the four-manifold $Q$ built by Cappell-Shaneson [10] reveals that the compact four-manifold $Z\mathrel{\mathop{:}}=Q\setminus\nu(\alpha_{Q})$ (where $\alpha_{Q}\subset Q$ is a simple loop that generates $\pi_{1}(Q)=\mathbb{Z}/2$) is homeomorphic, but not diffeomorphic to $D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2}$. The second clause of Theorem 32 encapsules work of Gompf, where he showed that the orientation two-cover of $Z$ is diffeomorphic to $D^{2}\times S^{2}$. The third clause of Theorem 32 follows from the property $Z\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}=(D^{2}\mathbin{\widetilde{\smash{\times}}}{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{2})\mathbin{\#}{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}$ [1, Section 3].
@@ 6.3. A spectral invariant as a diffeomorphism obstruction
$\bullet$ If the $\eta$-invariant of two closed nonorientable $\operatorname{Pin}^{+}$- four-manifolds $(Y,\phi_{Y})$ and $(Y^{\prime},\phi_{Y^{\prime}})$ with $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2=H^{1}(Y^{\prime};\mathbb{Z}/2)$ satisfies $\eta(Y,\phi_{Y})\neq\eta(Y^{\prime},\phi_{Y^{\prime}})\mod 2\mathbb{Z}$, then $Y_{1}$ is not stably diffeomorphic to $Y_{2}$.
@@ 6.4. $\operatorname{Pin}^{+}$-cobordism classes and a computational tool
$\bullet$ Kirby-Taylor [28, Theorem 5.2]. There is a group isomorphism
@@
$({\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4},\pm\phi_{{\mathbb{R}\mkern-0.5mu\mathrm{P}}^{4}})]=\pm 1\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
@@
$(Q,\pm\phi_{Q})]=\pm 9\in\Omega^{\operatorname{Pin}^{+}}_{4}\cong\mathbb{Z}/16.$
@@
Let $Y$ be a closed four-manifold with first cohomology group $H^{1}(Y;\mathbb{Z}/2)=\mathbb{Z}/2$ and that admits a $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$. Let $\alpha_{Y}\subset Y$ be an orientation-reversing simple loop such that $\alpha_{Y}]\neq 0\in H_{1}(Y;\mathbb{Z})$ and build
@@
[Proof]We first assemble the $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$ by gluing together $\operatorname{Pin}^{+}$-structures $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ and $(N_{p},\phi_{N_{p}})$ that induce the same $\operatorname{Pin}^{+}$-structure on their common boundary; see (6.14). The $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ induces a $\operatorname{Pin}^{+}$-structure $(Y\setminus\nu(\alpha_{Y}),\phi_{Y}^{0})$ as a codimension zero submanifold [28, Construction 1.5], which restricts to a $\operatorname{Pin}^{+}$-structure
@@
there are two choices of such structures [41, Corollary 6.4] and there is a one-to-one correspondence between the initial $\operatorname{Pin}^{+}$-structure $(Y,\phi_{Y})$ and the $\operatorname{Pin}^{+}$-structure (6.17) (see [27, Proposition 1 Chapter IV]). We now argue that there is a $\operatorname{Pin}^{+}$-structure $(N_{p},\phi_{N_{p}})$ that induces the $\operatorname{Pin}^{+}$-structure (6.17) on $\partial N_{p}=S^{2}\mathbin{\widetilde{\smash{\times}}}S^{1}$. Since $w_{2}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1})=0$ and $H^{1}(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1};\mathbb{Z}/2)=\mathbb{Z}/2$, there are two $\operatorname{Pin}^{+}$-structures on $S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1}$ [41, Corollary 6.4]. The hypothesis $p\neq 0\mod 2$ allows us to pick the $\operatorname{Pin}^{+}$-structure $(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\phi)$ that induces $(N_{p},\phi_{N_{p}})$ such that the induced $\operatorname{Pin}^{+}$-structure $(\partial N_{p},\phi_{N_{p}}|_{\partial})$ matches (6.17). We conclude that there is a $\operatorname{Pin}^{+}$-structure $(Y(p),\phi_{Y(p)})$. Next, we put together the compact five-manifold $V$ by gluing the product cobordisms
@@
$(S^{3}\mathbin{\widetilde{\smash{\times}}}S^{1},\pm\phi)]=0\in\Omega^{\operatorname{Pin}^{+}}_{4}$,